Golden Wedding Dress Company designs custom wedding dresses for brides to be. The person preparing the adjusting entries at year-end was unable to complete the adjustments due to illness. You have been given the following unadjusted trial balance along with some additional information for the December 31, 2017, year-end Unadjusted Account Balance $82,400 132,000 348,000 232,000 700 447,000 88,700 651,000 4,800 33,950 137,000 73,000 231,000 223,750 166,000 1,177,000 1,361,000 9,200 Accounts receivable Accum. deprec., building Accum. deprec., equipment Advance sales Allowance for doubtful accounts Building Cash Equipment Estimated warranty liability Income tax expense Land Merchandise inventory Mortgage payable Sarah Golden, capital Note payable Other operating expenses Sales Sales returns and allowances Other information 1. Assume all accounts have a normal balance 2.75% of the balance in the Advance Sales account is for wedding dresses to be made and delivered by Golden during 2018; the remaining 25% is from sales earned during 2017 3, Golden warranties its wedding dresses against defects and estimates its warranty liability to be 2% of adjusted net sales 4. The 3%, 5-year note payable was issued on October 1, 2017; interest is payable annually each September 30 5. A partial amortization schedule for the mortgage follows Year 2015 2016 2017 2018 2019 Interest Expense $11,986 11,106 10,192 9,240 8,250 Principal Portion $ 21,993 22,873 23,787 24,739 25,729 Annual Payment* $ 33,979 33,979 33,979 33,979 33,979 Principal Balance at Dec. 31 277,660 254,788 231,000 206,261 180,532 Payments are made annually each January 2 6. Uncollectible accounts are estimated to be 1.5% of outstanding receivables 7. A physical count of the inventory showed a balance actually on hand of $63,200 8, The balance in Income Tax Expense represents taxes accrued and paid for the 2017 year at the rate of $3,086 per month. Assume the income tax rate is 20%6, uncollectible accounts are estimated to be 1.5% of outstanding receivables. 7. A physical count of the inventory showed a balance actually on hand of $63,200. 8. The balance in Income Tax Expense represents taxes accrued and paid for the 2017 year at the rate of $3,086 per month. Assume the income tax rate is 20% Required 1. Based on the information provided, journalize the adjusting entries at December 31, 2017. (Round the final answers to 2 decimal places.) View transaction list Record to adjust for earned sales 2 Record the estimated warranty liability Record the accrual of interest expense on the note 1 3 payable. Record the accrual of interest on mortgage payable. Record to adjust for estimated uncollectible accounts. Record to adjust for shrinkage. Record the adjustment for income taxes owing. 4 bit Credit 5 6 7 Note:journal entry has been entered Record entry Clear entry View general journalGOLDEN WEDDING DRESS COMPANY Balance Sheet December 31, 2017 Assets Current assets: Total current assets Property, plant and equipment: Total property, plant and equipment Total assets 0 Liabilities Current liabilities: Total current liabilities 0 Non-current liabilities: Total liabilities Equity Total liabilities and equity
Answer #1

I hope this will help you.If requires clarification , you may comment below.

  1. Adjusting entries
  2. Date

    • Particulars
    • Debit
    • Credit

    December 31, 2017

    Advance sales

    $58,000

    Sales

    58,000

    (Entry recorded for adjusting earned sales of 25%)

    December 31, 2017

    Warranty expenses

    28,196

    Estimated Warranty liability

    28,196

    (Being entry recorded for warranty estimates = (1,361,000 + 58,000 – 9,200)*2% = 28,196)

    December 31, 2017

    Interest expense

    1,245

    Interest payable

    1,245

    (Being entry recorded for notes payable interest accrued from 1st Oct 2017 to Dec-17 = 166,000*3%*3/12 = 1245)

    December 31, 2017

    Interest expense

    10,192

    Interest payable

    10,192

    (Being Interest on mortgage has been recorded)

    December 31, 2017

    Allowance for doubt full debts

    1236

    Accounts receivable

    1236

    (Being 1.5% on accounts receivables classified as uncollectible = 82,400*1.5%=1,236)

    December 31, 2017

    Shrinkage expenses

    9,800

    Merchandised inventory

    9,800

    (Being shrinkage has been recorded = $73,000 – 63,200 = 9,800)

    December 31, 2017

    Income tax expense

    3,082

    Income tax payable

    3,082

    (Being Income tax on every month $3,086 , So =3086*12 =37032-33950)